# Quantum Curriculum Learning

### v3

@@ Abstract
We present a benchmark of Q-CurL for learning the approximation of the unitary dynamics of the spin-1/2 XY model with the Hamiltonian $H_{XY}=\sum_{j=1}^{Q}\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+h_{j}\sigma_{j}^{z}\right)$, where $h_{j}\in{\mathbb{R}}$ and $\sigma_{j}^{x},\sigma_{j}^{y},\sigma_{j}^{z}$ are the Pauli operators acting on qubit $j$. This model is important in the study of quantum many-body physics, as it provides insights into quantum phase transitions and the behavior of correlated quantum systems.
@@
To create the main task ${\mathcal{T}}_{M}$ and auxiliary tasks, we represent the time evolution of $H_{XY}$ via the ansatz $V_{XY}$, which is similar to the Trotterized version of $\exp(-i\tau H_{XY})$ [12]. The target unitary for the main task, $V^{(M)}_{XY}=\prod_{l=1}^{L_{M}}V^{(l)}(\bm{\beta}_{l})\prod_{l=1}^{L_{F}}V^{(l)}_{\textrm{fixed}}$, consists of $L_{M}=20$ repeating layers, where each layer $V^{(l)}(\bm{\beta}_{l})$ includes parameterized z-rotations RZ (with assigned parameter $\bm{\beta}_{l}$) and non-parameterized nearest-neighbor $\sqrt{i\textup{SWAP}}=\exp(\frac{i\pi}{8}(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}))$ gates. Additionally, we include the fixed-depth unitary $\prod_{l=1}^{L_{F}}V^{(l)}_{\textrm{fixed}}$ with $L_{F}=20$ layers at the end of the circuit $V^{(l)}$ to increase expressivity. Similarity, keeping the same ${\bm{\beta}}_{l}$, we create the target unitary for the auxiliary tasks ${\mathcal{T}}_{m}$ as $V^{(m)}_{XY}=\prod_{l=1}^{L_{m}}V^{(l)}({\bm{\beta}}_{l})\prod_{l=1}^{L_{F}}V^{(l)}_{\textrm{fixed}}$, with $L_{m}=1,2,\ldots,19$.
Figure 2(a) depicts the average HS distance over 100 trials of ${\bm{\beta}}_{l}$ and $V^{(l)}_{\textrm{fixed}}$ between the target unitary of each auxiliary task ${\mathcal{T}}_{m}$ (with $L_{m}$ layers) and the main task ${\mathcal{T}}_{M}$. We also plot the curriculum weight $c_{M,m}$ in Fig. 2(a) calculated in Eq. (5). Here, we consider the unitary $V_{XY}$ learning with $Q=4$ qubits via the hardware efficient ansatz $U_{\textrm{HEA}}({\bm{\theta}})$  [24, 23] and use $N=20$ Haar random states for input data ${\bm{x}}_{i}^{(m)}$ in each task ${\mathcal{T}}_{m}$. As depicted in Fig. 2(a), $c_{M,m}$ can capture the similarity between two tasks, as higher weights imply smaller HS distances.
Next, we propose a Q-CurL game to further examine the effect of Q-CurL. In this game, Alice has an ML model ${\mathcal{M}}({\bm{\theta}})$ to solve the main task ${\mathcal{T}}_{M}$, but she needs to solve all the auxiliary tasks ${\mathcal{T}}_{1},\ldots,{\mathcal{T}}_{M-1}$ first. We assume the data forgetting in task transfer, meaning that after solving task $A$, only the trained parameters ${\bm{\theta}}_{A}$ are transferred as the initial parameters for task $B$. We propose the following greedy algorithm to decide the curriculum order ${\mathcal{T}}_{i_{1}}\to{\mathcal{T}}_{i_{2}}\to\ldots\to{\mathcal{T}}_{i_{M}=M}$ before training. Starting ${\mathcal{T}}_{i_{M}}$, we find the auxiliary task ${\mathcal{T}}_{i_{M-1}}$ ($i_{M-1}\in\{1,2,\ldots,M-1\}$) with the highest curriculum weights $c_{i_{M},i_{M-1}}$. Similarity, to solve ${\mathcal{T}}_{i_{M-1}}$, we find the corresponding auxiliary task ${\mathcal{T}}_{i_{M-2}}$ in the remaining tasks with the highest $c_{i_{M-1},i_{M-2}}$, and so on. Here, curriculum weights $c_{i_{k},i_{k-1}}$ are calculated similarly to Eq. (5).
@@
Data-based Q-CurL.— We present a form of data-based Q-CurL that dynamically predicts the easiness of each sample at each training epoch, such that easy samples are emphasized with large weights during the early stages of training and conversely. Remarkably, it does not involve pre-training or additional training data, thereby avoiding any increase in quantum resource requirements.
Apart from improving generalization, data-based Q-CurL offers resistance to noise. This feature is particularly valuable in QML, where clean annotated data are often costly while noisy data are abundant. Existing QML models can accurately fit corrupted labels in the training data but often fail on test data [25]. We demonstrate that data-based Q-CurL enhances robustness by dynamically weighting the difficulty of fitting corrupted labels.
Figure 3: The test loss and accuracy of the trained QCNN (with and without using the data-based Q-CurL) in the quantum phase recognition task with 8 qubits under varying noise levels in corrupted labels. Here, the average and the best performance over 50 trials are plotted.
Inspired by the confidence-aware techniques in classical ML [19, 20, 21], the idea is to modify the empirical risk as
@@
Here, ${\bm{w}}=(w_{1},\ldots,w_{N})$, $\ell_{i}=\ell(h({\bm{x}}_{i}),{\bm{y}}_{i})$, and $w^{2}_{i}$ is the regularization term controlled by the hyper-parameter $\gamma>0$. The threshold $\eta$ distinguishes easy and hard samples with $e^{w_{i}}$ emphasizing the loss $l_{i}\ll\eta$ (easy sample) and neglecting the loss $l_{i}\gg\eta$ (hard samples, such as data with corrupted labels) [^1] In Supplementary Material, we have also discussed an interesting scenario where the modified loss in Eq. (7) can be used to emphasize complex quantum data during training, potentially reducing generation errors in quantum phase detection tasks under specific conditions. This aligns with the numerical results reported in Ref. [26], which appeared on arXiv after our paper.. The optimization is reduced to $\textup{min}_{{\bm{\theta}}}\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}})$, where ${\bm{\theta}}$ is the parameter of the hypothesis $h$. Here, $\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}})$ is decomposed at each loss $\ell_{i}$ and solved without quantum resources as $w_{i}=\textup{argmin}_{w}(l_{i}-\eta)e^{w}+\gamma w^{2}$. To control the difficulty of the samples, in each training epoch, we set $\eta$ as the average value of all $\ell_{i}$ obtained from the previous epoch. Therefore, $\eta$ adjusts dynamically in the early training stages but stabilizes near convergence.
@@
We apply the data-based Q-CurL to the quantum phase recognition task investigated in Ref. [10] to demonstrate that it can improve the generalization of the learning model. Here, we consider a one-dimensional cluster Ising model with open boundary conditions, whose Hamiltonian with $Q$ qubits is given by $H=-\sum_{j=1}^{Q-2}\sigma^{z}_{j}\sigma^{x}_{j+1}\sigma^{z}_{j+2}-h_{1}\sum_{j=1}^{Q}\sigma^{x}_{j}-h_{2}\sum_{j=1}^{Q-1}\sigma^{x}_{j}\sigma^{x}_{j+1}.$ Depending on the coupling constants $(h_{1},h_{2})$, the ground state wave function of this Hamiltonian can exhibit multiple states of matter, such as the symmetry-protected topological phase, the paramagnetic state, and the anti-ferromagnetic state. We employ the quantum convolutional neural network (QCNN) model [10] with binary cross-entropy loss for training. Without Q-CurL, we use the conventional loss $\hat{R}(h)=(1/N)\sum_{i=1}^{N}\ell_{i}$ for the training and test phase. In data-based Q-CurL, we train the QCNN with the loss $\hat{R}(h,{\bm{w}})$ while using $\hat{R}(h)$ to evaluate the generalization on the test data set. We use 40 and 400 ground state wave functions for the training and test phases, respectively (see [23] for details).

### v4

@@ II.1.2 Unitary learning task and Q-CurL game
$C_{\textrm{HST}}({\bm{\theta}})=\dfrac{d+1}{d}{\mathbb{E}}_{\ket{\psi}\sim\textrm{Haar}_{n}}\left[1-|\braket{\psi}{V^{\dagger}U({\bm{\theta}})}{\psi}|^{2}\right].$ (19)
This suggests a QML-based approach to learn the target unitary $V$, where we can access a training data set consisting of input-output pairs of pure $Q$-qubit states ${\mathcal{D}}_{{\mathcal{Q}}}(N)=\{(\ket{\psi}_{j},V\ket{\psi}_{j})\}_{j=1}^{N}$ drawn from the distribution ${\mathcal{Q}}$. If we take ${\mathcal{Q}}$ as the Haar distribution, we can instead train using the empirical loss:
@@
Figure 2: (a) The XY ansatz, employed for constructing target unitaries in both the main and auxiliary tasks, is inspired by the XY model with periodic boundary conditions. It comprises single-qubit $RZ$ rotations applied to qubit indices $1,3,\ldots$, and nearest-neighbor $\sqrt{\text{iSWAP}}$ gates, arranged with periodic boundary conditions. The placement of $\sqrt{\text{iSWAP}}$ gates in each layer $l$ varies depending on whether $l$ is odd or even. (b) The hardware-efficient (HE) ansatz, used to represent the hypothesis in the unitary learning task, consists of single-qubit RY and RZ rotations applied to all qubit indices, combined with nearest-neighbor CNOT gates, arranged with periodic boundary conditions.
@@
We present a benchmark of Q-CurL for learning the approximation of the unitary dynamics of the spin-1/2 XY model with the following Hamiltonian (with the periodic boundary condition):
$H_{XY}=\sum_{j=1}^{Q}\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+h_{j}\sigma_{j}^{z}\right),$ (21)
@@
where $h_{j}\in{\mathbb{R}}$ and $\sigma_{j}^{x},\sigma_{j}^{y},\sigma_{j}^{z}$ are the Pauli operators acting on qubit $j$. This model is important in the study of quantum many-body physics, as it provides insights into quantum phase transitions and the behavior of correlated quantum systems.
To create the main task ${\mathcal{T}}_{M}$ and auxiliary tasks, we represent the time evolution of $H_{XY}$ via the XY ansatz $V_{XY}$, which is similar to the Trotterized version of $\exp(-i\tau H_{XY})$ [20]. The target unitary for the main task consisting of $L_{M}=20$ repeating layers is defined as
$V^{(M)}_{XY}=\prod_{l=1}^{L_{M}}V^{(l)}(\bm{\beta}_{l})\prod_{l=1}^{L_{F}}V^{(l)}_{\textrm{fixed}},$ (22)
@@
where each layer $V^{(l)}(\bm{\beta}_{l})$ includes parameterized z-rotations RZ (with assigned parameter $\bm{\beta}_{l}$) and non-parameterized nearest-neighbor $\sqrt{i\textup{SWAP}}=\exp(\frac{i\pi}{8}(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}))$ gates [Fig. 2(a)]. Here, $\bm{\beta}_{l}$ are initialized randomly from a uniform distribution over $[0,1]$. Additionally, we include the fixed-depth unitary $\prod_{l=1}^{L_{F}}V^{(l)}_{\textrm{fixed}}$ with $L_{F}=20$ layers at the end of the circuit $\prod_{l=1}^{L_{M}}V^{(l)}(\bm{\beta}_{l})$ to increase expressivity. Each $V^{(l)}_{\textrm{fixed}}$ shares the same ansatz structure as $V^{(l)}(\bm{\beta}_{l})$, but with fixed $RZ$ gate angles assigned from a uniform distribution over $[0,2\pi]$. The target unitary for the main task $\mathcal{T}_{M}$ includes $(L_{M}+L_{F})\times\lceil Q/2\rceil=80$ RZ gate angles, where $L_{M}=20$, $L_{F}=20$, and $Q=4$.
Similarity, we create the target unitary for the auxiliary tasks ${\mathcal{T}}_{m}$ as
$V^{(m)}_{XY}=\prod_{l=1}^{L_{m}}V^{(l)}(\bm{\beta}_{l})\prod_{l=1}^{L_{F}}V^{(l)}_{\textrm{fixed}},$ (23)
with $L_{m}=1,2,\ldots,L_{M}-1$. Therefore, the auxiliary task $\mathcal{T}_{m}$ consists of $L_{m}$ quantum circuits, $V^{(1)}(\bm{\beta}_{1}),\ldots,V^{(L_{m})}(\bm{\beta}_{L_{m}})$, which form a subset of the $L_{M}$ quantum circuits $V^{(1)}(\bm{\beta}_{1}),\ldots,V^{(L_{M})}(\bm{\beta}_{L_{M}})$ in the main task ${\mathcal{T}}_{M}$. The parameter range for $\bm{\beta}_{l}$ is set to $[0,1.0]$ instead of $[0,2\pi]$ to ensure that the auxiliary tasks maintain a periodic similarity to the main task.
@@
Figure 4: The distribution and density of the training cost and test cost of the main task in the Q-CurL game, considering both random order and Q-CurL order based on the curriculum weights. Here, $N=20$ random input data are trained for 20 epochs with 100 trials of initial parameters in the model, and $N=20$ data are tested for each trained model. We consider two types of random input as (a) Haar-random $Q$-qubit states and (b) products of $Q$ Haar-random single-qubit states.
In our experiments, we consider the unitary learning with $Q=4$ qubits via the hardware efficient (HE) ansatz $U_{\textrm{HEA}}({\bm{\theta}})$. This ansatz comprises multiple blocks, where each block consists of single-qubit operations spanned by SU(2) on all qubits and two-qubit controlled-X entangling gates [46] repeated for all pairs of neighbor qubits. Here, we use rotation operators of Pauli Y and Z as single qubit gates. Mathematically, $U_{\textrm{HEA}}({\bm{\theta}})$ is defined as follows:
$~{}U_{\textrm{HEA}}({\bm{\theta}})=\prod_{l=1}^{L_{E}}\left(\prod_{q=1}^{Q}\left[U_{R}^{q,l}({\bm{\theta}}^{(l)})\right]\times U_{\text{Ent}}\right),$ (24)
with $Q$ qubits consisting of $L_{E}$ entangling gates $U_{\text{Ent}}$ alternating with rotation gates on each qubit. Here, we use $U_{R}({\bm{\theta}})=R_{Y}({\bm{\theta}}_{1})R_{Z}({\bm{\theta}}_{2})$, and $U_{\text{Ent}}$ is composed of CNOT gates placed in linear with indexes $(q,q+1)$ of qubits, arranged with periodic boundary conditions [Fig. 2(b)]. The number of parameters in this circuit is $2QL_{E}$. We configure the HE ansatz with $L_{E}=40$ layers to achieve high expressivity, resulting in 320 parameters for $Q=4$ qubits. These parameters are initialized randomly from a uniform distribution over $[0,2\pi]$.
Figure 3(a) depicts the average HS distance over 100 trials of ${\bm{\beta}}_{l}$ and $V^{(l)}_{\textrm{fixed}}$ between the target unitary of each auxiliary task ${\mathcal{T}}_{m}$ (with $L_{m}$ layers) and the main task ${\mathcal{T}}_{M}$. The Hilbert-Schmidt (HS) distance $\Delta_{M,m}$ represents the distance between the unitaries $V^{(m)}_{XY}$ and $V^{(M)}_{XY}$. It is calculated as follows, assuming the explicit forms of $V^{(m)}_{XY}$ and $V^{(M)}_{XY}$ are known:
[…]
where $d=2^{Q}$ denotes the dimension of the Hilbert space. The HS distance here depends solely on the design of the target unitaries for the auxiliary tasks and the main task, and is independent of the trainable ansatz $U_{\textup{HEA}}(\bm{\theta})$. Thus, it offers insight into which auxiliary task is most similar to the main task by comparing the distances $\Delta_{M,1},\Delta_{M,2},\ldots,\Delta_{M,M-1}$. As shown in Fig. 3(a), the HS distances at auxiliary tasks $\mathcal{T}_{8}$ and $\mathcal{T}_{19}$ are approximately 0.6, which are considerably smaller than those at other auxiliary tasks (approximately 1.0). As the HS distance cannot be computed in advance, we rely on the curriculum weight to guide the design of the curriculum.
We plot the curriculum weight $c_{M,m}$ in Fig. 3(a) calculated in Eq. (18) with the basis functions form in Eq. (9). Here, we use $N=20$ Haar random states for input data ${\bm{x}}_{i}^{(m)}$ in each task ${\mathcal{T}}_{m}$. As depicted in Fig. 3(a), $c_{M,m}$ can capture the similarity between two tasks, as higher weights imply smaller HS distances.
Next, we propose a Q-CurL game to further examine the effect of Q-CurL. In this game, Alice has an ML model ${\mathcal{M}}({\bm{\theta}})$ to solve the main task ${\mathcal{T}}_{M}$, but she needs to solve all the auxiliary tasks ${\mathcal{T}}_{1},\ldots,{\mathcal{T}}_{M-1}$ first. We assume the data forgetting in task transfer, meaning that after solving task $A$, only the trained parameters derived from task $A$ are transferred as the initial parameters for task $B$. The Q-CurL framework provides an algorithm to determine an efficient order of auxiliary tasks to facilitate solving the main task. We propose the following greedy algorithm to decide the curriculum order ${\mathcal{T}}_{i_{1}}\to{\mathcal{T}}_{i_{2}}\to\ldots\to{\mathcal{T}}_{i_{M}=M}$ before training. Starting ${\mathcal{T}}_{i_{M}}$, we find the auxiliary task ${\mathcal{T}}_{i_{M-1}}$ ($i_{M-1}\in\{1,2,\ldots,M-1\}$) with the highest curriculum weights $c_{i_{M},i_{M-1}}$. Similarity, to solve ${\mathcal{T}}_{i_{M-1}}$, we find the corresponding auxiliary task ${\mathcal{T}}_{i_{M-2}}$ in the remaining tasks with the highest $c_{i_{M-1},i_{M-2}}$, and so on. Here, curriculum weights $c_{i_{k},i_{k-1}}$ are calculated similarly to Eq. (18).
@@
Figure 4 depicts the distribution and density of the train loss and test loss of the main task in the Q-CurL game, comparing the Q-CurL order with a random order. Here, $N=20$ random input data are trained for 20 epochs with 100 trials of initial parameters in the model, and $N=20$ data are tested for each trained model. We consider two types of random inputs as (a) Haar-random $Q$-qubit states [Fig. 4(a)] and (b) products of $Q$ Haar-random single-qubit states [Fig. 4(b)]. In both types of input states, the order in solving the Q-CurL game derived via the task-based Q-CurL method outperforms the performance when considering the random order.
The Q-CurL game setting and the heuristic greedy algorithm discussed here demonstrate the usefulness of using curriculum weight to decide the curriculum order. We can further explore several variations of the Q-CurL game. For instance, instead of using the test loss ${\mathcal{L}}^{(M)}_{t}$ of the main task ${\mathcal{T}}_{M}$ as the evaluation metric for the curriculum order ${\mathcal{T}}_{i_{1}}\to{\mathcal{T}}_{i_{2}}\to\ldots\to{\mathcal{T}}_{i_{M}=M}$, one could consider minimizing the total test loss $\sum_{k=2}^{M}{\mathcal{L}}^{(i_{k})}_{t}$ This approach would lead to a heuristic algorithm aimed at maximizing the total curriculum weights $\sum_{k=2}^{M}c_{i_{k},i_{k-1}}$. Another variation is to consider the task difficulty perspective. For example, we could set the first task to be solved initially (as we know it is easy to solve, or we already have a trained model) and then determine an optimal task order that smoothly transitions from the first task to the main task.
II.2 Data-based Q-CurL
We present a form of data-based Q-CurL that dynamically predicts the easiness of each sample at each training epoch, such that easy samples are emphasized with large weights during the early stages of training and conversely. Remarkably, it does not involve pre-training or additional training data, thereby avoiding any increase in quantum resource requirements.
Apart from improving generalization, data-based Q-CurL offers resistance to noise. This feature is particularly valuable in QML, where clean annotated data are often costly while noisy data are abundant. Existing QML models can accurately fit corrupted labels in the training data but often fail on test data [47]. We demonstrate int the quantum phase recognition task that data-based Q-CurL enhances robustness by dynamically weighting the difficulty of fitting corrupted labels.
II.2.1 Dynamical learning schedule
In the procedure without using the Q-CurL, we use the conventional loss as the empirical risk $\hat{R}(h)=\dfrac{1}{N}\sum_{i=1}^{N}\ell_{i}$ for the training and testing phase. In data-based Q-CurL, inspired by the confidence-aware techniques in classical ML [38, 35, 36], we modify the conventional loss to the following form of the dynamical loss function:
@@ II.2.1 Dynamical learning schedule
Here, ${\bm{w}}=(w_{1},\ldots,w_{N})$ and $w^{2}_{i}$ is the regularization term controlled by the hyper-parameter. The threshold $\eta$ distinguishes easy and hard samples with $e^{w_{i}}$ emphasizing the loss $\ell_{i}\ll\eta$ (easy sample) or the loss $\ell_{i}\gg\eta$ (hard samples, such as data with corrupted labels). The optimization is reduced to
$\textup{min}_{{\bm{\theta}}}\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}}),$ (27)
where ${\bm{\theta}}$ is the parameter of the hypothesis $h$. Here, $\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}})$ is decomposed at each loss $\ell_{i}$ and solved without quantum resources as
$w_{i}=\textup{argmin}_{w}(l_{i}-\eta)e^{w}+\gamma w^{2}.$ (28)
To control the difficulty of the samples, in each training epoch, we set $\eta$ as the average value of all $\ell_{i}$ obtained from the previous epoch. Therefore, $\eta$ adjusts dynamically in the early training stages but stabilizes near convergence.
In Appendix B, we present the details of solving Eq. (28). Given the solution of Eq. (28), by controlling the sign of $\gamma$, the dynamical loss can be used to prioritize emphasizing the small losses (with $\gamma>0$) or the large losses (with $\gamma<0$). We define these two scenarios as easy Q-CurL and hard Q-CurL, respectively.
The easy Q-CurL aligns with the classical curriculum learning context [36]. This approach is particularly beneficial when hard samples, such as those with noisy labels, could mislead the optimization process. However, the hard Q-CurL can be advantageous in scenarios where hard samples such as the complex quantum data are crucial for guiding the model to extract essential features without being distracted by irrelevant ones. We will provide an example of this interesting scenario at the end of the next subsection.
Figure 5: The schematic diagram for the QCNN [18] used in our quantum phase recognition task. Here, C, P, and FC represent convolutional, pooling, and fully connected layers, respectively. In the convolutional layer, local unitaries $V_{kl}$ are applied to pairs of neighboring qubits $(k,l)$ in the illustrated order, excluding those previously discarded, under periodic boundary conditions. All $V_{kl}$ in the same layer share the same parameters. In the pooling layer, qubits with even indices among the remaining qubits are discarded. This sequence of alternating convolutional and pooling layers ends with a fully connected layer, which operates as a single convolutional operator on the remaining qubits. Finally, we apply the Hadamard gate to the last remaining qubit and then perform a measurement in the Z-basis to classify the input quantum data $\rho_{\textup{in}}$.
Figure 6: (a) The test loss and accuracy of the trained QCNN (with and without using the data-based Q-CurL) in the quantum phase recognition task with 8 qubits under varying noise levels in corrupted labels. Here, the average and the best performance over 50 trials are plotted. (b) The test loss (left panel) and test accuracy (right panel) of the trained QCNN on the quantum phase recognition task with (solid lines) or without (dotted lines) using the data-based Q-CurL over different numbers of qubits. Here, we consider two different noise levels in the corrupted training labels: $p=0.2$ (blue) and $p=0.3$ (red).
II.2.2 Quantum phase recognition task
@@
We apply the data-based Q-CurL to the quantum phase recognition task investigated in Ref. [18] to demonstrate that it can improve the generalization of the learning model. Here, we consider a one-dimensional cluster Ising model with open boundary conditions, whose Hamiltonian with $Q$ qubits is given by
$~{}H=-\sum_{i=1}^{Q-2}\sigma^{z}_{i}\sigma^{x}_{i+1}\sigma^{z}_{i+2}-h_{1}\sum_{i=1}^{Q}\sigma^{x}_{i}-h_{2}\sum_{i=1}^{Q-1}\sigma^{x}_{i}\sigma^{x}_{i+1}.$ (29)
Depending on the coupling constants $(h_{1},h_{2})$, the ground state wave function of this Hamiltonian can exhibit multiple states of matter, such as the symmetry-protected topological phase (SPT phase), the paramagnetic state, and the anti-ferromagnetic state.
We employ the quantum convolutional neural network (QCNN) model [18] to determine the matter phase of quantum states. Inspired by classical convolutional neural networks, the QCNN model consists of convolutional, pooling, and fully connected layers. The convolutional layers use local unitary gates to extract local features from the input data, while the pooling layers reduce the number of qubits. This alternation of layers ends in a fully connected layer that functions as a single convolution operator on the remaining qubits, providing an output through the measurement of the final qubit. The QCNN is governed by variational parameters that are optimized to classify training data accurately. In our implementation, the convolutional and fully connected layers are constructed using the Pauli decomposition of two-qubit unitary gates $V_{kl}$, expressed as $V_{kl}=\prod_{j=1}^{15}e^{-i\theta_{j}P_{j}}$ with 15 parameters, where $\{P_{j}\}$ are the Pauli operators for two qubits, excluding the identity matrix. Here, $V_{kl}$ applies to pairs of neighboring qubits $(k,l)$ with the order illustrated in Fig. 5, excluding those previously discarded. Each layer utilizes the same parameters for all unitary gates. In the pooling layer, qubits with even indices among the remaining qubits are discarded. Before measuring the output, we apply the Hadamard gate to the remaining qubit and then perform a measurement in the Z-basis. In this QCNN scheme, the number of independent parameters is given by $15\times\lceil\log_{2}(Q)\rceil$, where $\lceil\log_{2}(Q)\rceil$ represents the number of convolutional and fully connected layers.
For each training quantum data |$\ket{\psi_{i}}$ and its corresponding label $y_{i}$, the QCNN produces the output $q_{i}$ ($-1\leq q_{i}\leq 1$). The single loss $\ell_{i}$ is defined using the binary cross-entropy (BCE) loss as follows:
$\ell_{i}=-y_{i}\log(\hat{y}_{i})-(1.0-y_{i})\log(1.0-\hat{y}_{i}),$ (30)
where $\hat{y}_{i}=\textup{sigmoid}(\mu q_{i})$. Here, we consider the scaling output with the coefficient $\mu=1.0$. The label is predicted as 00 if $\hat{y}_{i}<0.5$ and $1$ if $\hat{y}_{i}\geq 0.5$. In the procedure without using the Q-CurL, we use the conventional loss $\hat{R}(h)=\dfrac{1}{N}\sum_{i=1}^{N}\ell_{i}$ for the training. We also use $\hat{R}(h)$ to evaluate the generalization on the test data set.
Similar to the setup in Ref. [18], we generate a training set of 40 ground state wave functions corresponding to $h_{2}=0$ and $h_{1}$ sampled at equal intervals in [0.0, 1.6]. The state is analytically solvable for these parameter choices, and this solution is used to label the training dataset (0 for the paramagnetic or antiferromagnetic phase and 1 for the SPT phase). The ground truth phase boundaries, which separate the two phases, are determined using DMRG simulations. Based on these boundaries, we also create a test dataset of 400 ground state wave functions corresponding to $h_{2}\in$0.8439, 0.6636, 0.5033, 0.3631, 0.2229, 0.09766, -0.02755, -0.1377, -0.2479, -0.3531}, and $h_{1}$ sampled 40 times at equal intervals in [0.0, 1.6]. The optimization is performed by the Adam method with a learning rate of 0.001 and 500 epochs of training.
In our experiment, we consider the scenario of fitting corrupted labels. Given a probability $p$ ($0\leq p\leq 1$) representing the noise level, the true label $y_{i}\in\{0,1\}$ of quantum state |$\ket{\psi_{i}}$ is transformed to the corrupted label $1-y_{i}$ with probability $p$, while it remains the true label with probability $1-p$.